OpenStudy (anonymous):

Find the cube roots of 27(cos 279° + i sin 279°).

OpenStudy (anonymous):

27(cos 279° + i sin 279°) because it is a cube root we divide 360°/3 = 120° so we add 120° to every root and we need to find the third root z1 = 27(cos 279° + i sin 279°) z2 = 27(cos 39° + i sin 39°) z3 = 27(cos 159° + i sin 159°) (27 * cos 159°) + (27 * i sin 159°) -25.20667152 + 9.675934638i

OpenStudy (anonymous):

\[Let z=r \left( \cos \theta+\iota \sin \theta \right)=re ^{\iota \theta}=re ^{\iota \left( 2n \pi+\theta \right)} \] \[z ^{\frac{ 1 }{ 3 }}=r ^{\frac{ 1 }{3 }}e ^{\frac{ \iota \left( 2n \pi+\theta \right) }{ 3 }} \] \[z ^{\frac{ 1 }{3 }}=r ^{\frac{ 1 }{ 3 }}\left\{ \cos \frac{ 2n \pi+\theta }{ 3 } +\iota \sin \frac{ 2n \pi+\theta }{ 3 }\right\}\] plug r=27,n=0,1,2, \[\theta=279 \] solve and get three solutions.